# Advances in Radiance Field for Dynamic Scene: From Neural Field to Gaussian Field

### v1

@@ 3.1.1 Rigid Motion
Rigid motion encompasses transformations in which an object preserves its shape and size while undergoing rotation and translation. Internally, distances and angles remain unchanged, and thus no deformation occurs. This can be mathematically expressed via a rotation matrix $\boldsymbol{R}\in SO(3)$ (or quaternion $\boldsymbol{q}\in\mathbb{R}^{4}$) and a translation vector $\boldsymbol{t}\in\mathbb{R}^{3}$, such that:
$\mathbf{x}_{t}\;=\;\mathcal{T}_{\theta}\bigl{(}\mathbf{x}_{t-1}\bigr{)}\;=\;\boldsymbol{R}\,\mathbf{x}_{t-1}\;+\;\boldsymbol{t}.$ (7)
@@ 3.1.2 Articulated Motion
Articulated motion, also known as piecewise rigid motion, describes a class of transformations where individual segments (e.g., limbs in a skeletal model) undergo rigid transformations while the object’s overall motion appears non-rigid due to the relative movement between segments. This type of motion is ubiquitous in both natural organisms (human and animal skeletons) and manufactured systems (robotic arms, mechanical assemblies, and hinged objects).
@@
An articulated object is typically represented by a kinematic tree–a hierarchical structure of rigid segments connected by joints that provide specific degrees of freedom (e.g., rotation or translation). This hierarchical organization captures how local transformations at each joint propagate through the kinematic chain. The human body represents a quintessential example of such articulated structures, and various parameterization strategies, such as SCAPE [81], SMPL [33], SMPL-X [82], and MANO [34], have produced influential frameworks for modeling its complex motion. Among these, the Skinned Multi-Person Linear Model (SMPL) has gained particular prominence due to its balance of expressiveness and computational efficiency.
SMPL conceptualizes the human body as a kinematic tree with 24 joints, each undergoing a rigid transformation relative to its parent, with the root joint defining the transformation from the body’s local space to world space. The model employs linear blend skinning (LBS) to deform a canonical rest pose according to a target pose configuration. By assigning blend weights to each vertex on a canonical mesh, SMPL effectively encodes how much each joint’s transformation influences that point. Formally, for a point $\mathbf{x}^{c}_{i}$ in the canonical space, its deformed position $\mathbf{x}^{p}_{i}$ in the posed space is computed by:
$\mathbf{x}^{p}_{i}\;=\;\mathcal{T}_{\theta}\!\bigl{(}\mathbf{x}^{c}_{i};\mathcal{w}(\mathbf{x}^{c}_{i}),\pi(J)\bigr{)}\;=\;\sum_{j=1}^{J}\bigl{(}w_{j}\!\bigl{(}\mathbf{x}^{c}_{i}\bigr{)}\,T_{j}\,\bigr{)}\mathbf{x}^{c}_{i},$ (8)
@@
where $\pi(J)$ represents the joint pose parameters, $w_{j}(\mathbf{x}^{c}_{i})$ are blend weights determining joint $j$’s influence on point $\mathbf{x}^{c}_{i}$, and $T_{j}\in\mathbb{R}^{4\times 4}$ is the rigid transformation matrix of the $j$-th joint. SMPL predefines these blend weights on a canonical mesh parametrized by a shape code $\beta$, allowing for personalized body proportions.
@@
While LBS accurately describes transformations for points on the template surface, handling points in free space (e.g., for volumetric rendering) requires additional techniques. Traditional approaches use barycentric interpolation or nearest-neighbor methods to extend blend weights to off-surface points [30, 36, 83]. However, these approaches may struggle with complex deformations, especially for loose clothing or accessories. To address these limitations, recent approaches employ neural skinning fields that replace or augment traditional pre-defined blend weights. A learnable skinning function maps a point $\mathbf{x}$ to its joint-influence weights:
${\mathcal{W}}(\mathbf{x})\;=\;{\mathcal{S}}_{\theta}\!\bigl{(}\mathbf{x},\,\pi(t)\bigr{)},$ (9)
@@
where $\pi(t)$ encodes time-dependent variations. Learning such fields from scratch presents challenges; therefore, many methods incorporate the pre-defined SMPL blend weights as a strong prior, allowing the learned skinning weights to deviate moderately from the canonical configuration when necessary [35, 84].
@@ 3.1.3 Non-Rigid Motion
Non-rigid motion permits objects to deform by locally changing the relative positions of points and is therefore indispensable for describing cloth folding, facial expression, fluid flow, and other complex dynamics. Unlike rigid or articulated motion, which can exploit kinematic templates or low-dimensional parametric models, general non-rigid deformation typically lacks structured representations due to potential topology changes and the infinite degrees of freedom involved. Classical tracking pipelines rely on sparse feature matching, yielding motion fields that are far too coarse to capture fine-scale deformation. Physically based methods may provide accurate solutions, yet their computational cost becomes prohibitive for large scenes or real-time use.
Recent advances have reformulated non-rigid motion as an implicit, learnable neural field . These approaches employ compact neural networks $\Delta_{\theta}$ to predict dense, continuous displacement fields of the form:
$\mathbf{x}_{t}\;=\;\mathbf{x}_{t-1}\;+\;\Delta_{\theta}\!\bigl{(}\mathbf{x}_{t-1};\pi(t)\bigr{)}.$ (10)
where $\Delta_{\theta}$ maps any 3D point at time $t-1$ to its offset at time $t$, conditioned on a temporal code $\pi(t)$. Such neural fields can be trained directly from 2D images through differentiable rendering, dispensing with explicit 3D supervision and delivering per-point motion estimates. The learning problem, however, is ill-posed—especially under monocular observations—because many distinct deformations can explain the same pixel sequence. Contemporary systems therefore narrow the solution space with a hierarchy of priors. Geometric regularizers promote smooth, continuous flows and discourage fold-overs or self-intersections [13]; isometry or near-rigidity assumptions penalize changes in pairwise distances [49, 85]; additional 2D cues such as dense optical flow or sparse feature correspondences supply direct motion evidence [19, 86]; and, when multi-view images are available, cross-view consistency enforced via neural rendering further suppresses spurious solutions [87]. A detailed discussion of these auxiliary cues and their implementation appears in Sec. 4.5.
@@ 3.1.4 Hybrid Motion
Most real-world scenes exhibit a mixture of motion types–rigid, articulated, and general non-rigid–whose simultaneous presence gives rise to hybrid motion. A prime example is the human body: a primarily articulated skeleton undergoes a global rigid transformation, while soft tissue, loose clothing, and hair introduce finer, highly deformable dynamics [88, 89]. Capturing this interplay demands a representation that is both structured enough to model global motion and flexible enough to accommodate local deviations.
@@
In this formulation, $\mathcal{T}_{\theta_{1}}$ represents a relatively constrained, interpretable global motion model (e.g., rigid transformation or articulated skeleton deformation), while $\Delta_{\theta_{2}}$ is typically implemented as a neural field that predicts residual displacements to capture complex, local deformations. Besides effective human and animal modeling, similar approaches have been applied to other domains with composite motion, including deformable objects with near-rigid parts and multi-object scenes with varying motion characteristics [80, 51]. This hierarchical decomposition offers several advantages: it maintains physical interpretability by isolating well-defined transformations; it reduces the complexity of the learning problem by having the neural component focus on residual details rather than the entire motion; and it provides explicit control over coarse motion while allowing data-driven capture of fine details such as clothing wrinkles, or facial expressions. Despite implementation challenges in properly balancing and integrating different motion representations, hybrid approaches remain essential for high-fidelity dynamic scene reconstruction, realistic animation, and applications requiring both structural accuracy and fine detail preservation.
Figure 5: Illustration of typical motion representation methods. An intuitive approach extends the 5D plenoptic function to 6D by incorporating time $t$ as an additional dimension, creating a 4D Spacetime (a) representation. Alternative methods represent the scene as a shared reference space and time-dependent Deformation Fields (b). This reference can be a globally shared Canonical Space (b-1) consistent across the entire video, or Keyframes (b-2) shared locally among neighboring frames. When each frame serves as the keyframe for its successive frame, the representation becomes a frame-to-frame Flow Field (c). Point Tracking (d) continuously integrates point trajectories across frames, rather than using discrete frame-to-frame correspondences. Factorization (e) approaches decompose 4D spacetime into Hyperplanes (e-1) or represent complex motion using low-rank Basis (e-2) functions, effectively reducing dimensionality while preserving motion characteristics.
@@ 3.2.1 Representing via 4D Spacetime
Starting from coordinate-based representations for static scenes [1], which utilize the 5D plenoptic function to represent light fields as $\mathcal{F}_{\theta}\colon(\mathbf{x},\mathbf{d})\longmapsto(\mathbf{c},\sigma)$, approaches for dynamic scenes naturally extend the input domain to 6D by incorporating temporal information $t$ [26, 90, 16, 40], resulting in $\mathcal{F}_{\theta}\colon(\mathbf{x},\mathbf{d},t)\longmapsto(\mathbf{c},\sigma)$. More generally, methods may condition on alternative temporal encodings beyond direct time input, such as frame indices $i$ or learned per-frame latent codes $\ell_{i}$ [91, 41, 92].
With this formulation, each frame of the dynamic scene is optimized independently through frame-by-frame optimization, leveraging multi-view observations to minimize a rendering loss between the synthesized output $I_{r}$ and the ground-truth image $I_{gt}$:
$\arg\min_{\theta}\sum\left\|I_{r}-I_{gt}\right\|_{2}^{2},\ where\ I_{r}=\mathcal{R}\Bigl{(}\mathcal{F}_{\theta}\bigl{(}\mathbf{x},\mathbf{d},\pi(t)\bigr{)}\Bigr{)},$ (12)
where $\mathcal{R}$ represents a differentiable volume rendering function (e.g., ray tracing), $\mathcal{F}_{\theta}$ denotes the radiance field with learnable parameters $\theta$, and $\pi(t)$ encodes temporal information. Motion is thus implicitly encoded within the same radiance field that represents the scene—rather than being handled by a dedicated motion module—and is supervised solely through available 2D image observations.
While this 4D spacetime representation with frame-by-frame optimization can produce high-fidelity reconstructions at individual time steps, it suffers from notable limitations. Using a single time-related variable to represent frame-level motion makes these models can not provides explicit point-level motion information. Furthermore, without an explicit 3D scene model, maintaining cross-view consistency between novel viewpoints becomes challenging. Despite these limitations, 4D spacetime representations remain central in many dynamic scene reconstruction pipelines, particularly when memory efficiency and 4D video compression are priorities.
3.2.2 Representing via Canonical Space with Deformation Field
A prevalent approach for modeling dynamic scenes decomposes the representation into a static canonical space and time-varying deformation fields. This canonical space–often designated as a "reference frame"–serves as a common coordinate system from which all observed frames are derived through learned deformations. For rigid or articulated objects, this canonical space typically corresponds to a rest pose (e.g., initial frame or neutral skeletal configuration). For non-rigid scenes, it should capture sufficient geometric detail to facilitate robust correspondence estimation across the sequence.
@@ 3.2.2 Representing via Canonical Space with Deformation Field
$\Delta_{i\to c}(\mathbf{x}_{i})\;=\;\Psi_{\theta}\bigl{(}\mathbf{x}_{i},\;\pi(t)\bigr{)}.$ (14)
@@
Within this framework, motion between any two observation frames $i$ and $j$ can be computed via the canonical space:
@@
For scenes with complex motions or topology changes where a single global canonical space becomes impractical, other approaches employ multiple local canonical spaces or keyframes that serve as reference frames for temporal subwindows [41, 95]. This extension enables the representation of longer sequences with more complex dynamics.
By decoupling the scene into a static reference space and time-dependent deformations, this framework accommodates diverse motion types ranging from rigid transformations to fully non-rigid deformations. It also facilitates transparent analysis and manipulation of motion—valuable in medical applications tracking tissue deformation and entertainment contexts requiring precise animation control. The explicit modeling of deformations further enables motion transfer, pose editing, and novel pose synthesis applications that are challenging with implicit spacetime representations.
@@ 3.2.3 Representing via Flow Field
A more direct strategy for modeling dynamic scenes involves capturing motion between consecutive frames rather than relating observations to a canonical space. This frame-to-frame approach leverages incremental deformations to represent complex motions, decomposing significant transformations into smaller, more tractable steps. Such representations naturally accommodate topology changes and extreme deformations that challenge canonical space methods.
Flow fields provide a principled framework for modeling these consecutive-frame dynamics. In their continuous form, velocity fields [96, 50, 19] specify instantaneous motion by assigning a velocity vector $\mathbf{v}(\mathbf{x},t)$ to each point $\mathbf{x}$ in space at time $t$. The trajectory of a point initially at position $\mathbf{x}_{t-1}$ can be computed by integrating this velocity field:
@@
where $\mathbf{v}(\mathbf{x}(\tau),\tau)$ represents the velocity at intermediate time $\tau$. While theoretically elegant, obtaining dense ground-truth velocities in practical settings is often infeasible.
For discrete time steps, scene flow fields $\Phi(\mathbf{x},t)$ directly model displacement vectors between consecutive frames:
$\Phi(\mathbf{x},t)\;=\;\mathbf{x}_{t}\;-\;\mathbf{x}_{t-1}.$ (17)
Scene flow and velocity fields are mathematically related—velocity represents the time derivative of scene flow—expressed differentially as:
$\frac{\partial\Phi}{\partial t}(\mathbf{x},t)\;=\;\mathbf{v}\bigl{(}\Phi(\mathbf{x},t),\,t\bigr{)},\quad\text{s.t.}\quad\Phi(\mathbf{x},\,t-1)=\mathbf{x}_{t-1},$ (18)
In practical implementations, both fields are typically parameterized by neural networks and optimized through differentiable rendering pipelines.
The frame-to-frame motion between frames $i$ and $j$ is expressed as:
$\Delta_{i\to j}\;=\;\Phi_{\theta}\bigl{(}\mathbf{x},\,\pi(t_{i})\bigr{)}.$ (19)
These local flow fields can be composed to represent extended trajectories through:
@@
This frame-to-frame paradigm offers two significant advantages. First, it operates on incrementally small displacements that neural networks can more effectively learn and generalize across diverse scenarios. Second, it eliminates the need for a global canonical space—often problematic under significant deformations or topology changes. These properties make flow-based representations particularly suitable for complex non-rigid dynamics where corresponding points may appear, disappear, or undergo substantial deformation throughout a sequence.

### v2

@@ 3.1.1 Rigid Motion
Rigid motion encompasses transformations in which an object preserves its shape and size while undergoing rotation and translation. Internally, distances and angles remain unchanged, and thus no deformation occurs. This can be mathematically expressed via a rotation matrix $\mathbf{R}\in SO(3)$ (or quaternion $\mathbf{q}\in\mathbb{R}^{4}$) and a translation vector $\mathbf{t}\in\mathbb{R}^{3}$:
$\mathbf{x}_{t}\;=\;\mathcal{T}_{\theta}\bigl{(}\mathbf{x}_{t-1}\bigr{)}\;=\;\mathbf{R}\,\mathbf{x}_{t-1}\;+\;\mathbf{t}.$ (5)
@@ 3.1.2 Articulated Motion
Articulated motion, also known as piecewise rigid motion, describes a class of transformations where individual segments (e.g., limbs in a skeletal model) undergo rigid transformations while the object’s overall motion appears non-rigid due to the relative movement between segments. This type of motion exists in both natural organisms (human and animal skeletons) and manufactured systems (robotic arms, mechanical assemblies, and hinged objects).
@@
An articulated object is typically represented by a kinematic tree–a hierarchical structure of rigid segments connected by joints that provide specific degrees of freedom (e.g., rotation or translation). This hierarchical organization captures how local transformations at each joint propagate through the kinematic chain. The human body represents a quintessential example of such articulated structures, and various parameterization strategies, such as SCAPE [73], SMPL [33], SMPL-X [74], and MANO [34], have produced influential frameworks for modeling its complex motion, among which SMPL has gained particular prominence due to its balance of expressiveness and computational efficiency. SMPL conceptualizes the human body as a kinematic tree with 24 joints, each undergoing a rigid transformation relative to its parent, with the root joint defining the transformation from body space to world space. The model employs linear blend skinning (LBS) to deform a canonical rest pose according to a target pose configuration. By assigning blend weights to each vertex on a canonical mesh, SMPL effectively encodes how much each joint’s transformation influences that point. Formally, for a point $\mathbf{x}^{c}_{i}$ in the canonical space, its deformed position $\mathbf{x}^{p}_{i}$ in the posed space is computed as follows:
$\mathbf{x}^{p}_{i}\;=\;\mathcal{T}_{\theta}\!\bigl{(}\mathbf{x}^{c}_{i},\mathbf{w}(\mathbf{x}^{c}_{i});\pi(J)\bigr{)}\;=\;\sum_{j=1}^{J}\bigl{(}w_{j}\!\bigl{(}\mathbf{x}^{c}_{i}\bigr{)}\,\textbf{T}_{j}\,\bigr{)}\mathbf{x}^{c}_{i},$ (6)
@@
where $\pi(J)$ represents the pose parameters, $w_{j}(\mathbf{x}^{c}_{i})$ are blend weights that determine the influence of $j$ th joint on point $\mathbf{x}^{c}_{i}$, and $\textbf{T}_{j}\in\mathbb{R}^{4\times 4}$ is the rigid transformation of the $j$-th joint.
@@
While LBS accurately describes transformations for points on the template surface, handling points in free space (e.g., for volumetric rendering) requires additional techniques. Traditional approaches use barycentric interpolation or nearest-neighbor methods to extend blend weights to points outside the surface [30, 36, 75]. However, these approaches may struggle with complex deformations, especially for loose clothing or accessories. To address these limitations, recent approaches employ neural skinning fields that replace or augment traditional pre-defined blend weights. A learnable skinning function maps a point $\mathbf{x}$ to its blending weights:
$\mathbf{w}(\mathbf{x})\;=\;{\mathcal{S}}_{\theta}\!\bigl{(}\mathbf{x};\pi(t)\bigr{)},$ (7)
@@
where $\pi(t)$ encodes time-dependent conditions, e.g. pose parameters. Learning such fields from scratch presents challenges; therefore, many methods incorporate the predefined SMPL blend weights as a strong prior, allowing the learned skinning weights to deviate moderately from the canonical configuration when necessary [35, 76].
@@ 3.1.3 Non-Rigid Motion
Non-rigid motion permits objects to deform by locally changing the relative positions of points and is therefore indispensable for describing cloth folding, facial expression, fluid flow, and other complex dynamics. Unlike rigid or articulated motion, which can exploit kinematic templates or low-dimensional parametric models, general non-rigid deformation typically lacks structured representations due to potential topology changes and the infinite degrees of freedom. Classical tracking pipelines rely on sparse feature matching, yielding motion fields that are far too coarse to capture fine-scale deformation. Physically based methods may provide accurate solutions, yet their computational cost becomes prohibitive for large scenes or real-time use.
Recent advances have reformulated non-rigid motion as an implicit, learnable neural field. These approaches employ compact neural networks to predict dense, continuous displacement fields of the form:
$\mathbf{x}_{t}=\mathcal{T}_{\theta}\bigl{(}\mathbf{x}_{t-1};\pi(t)\bigr{)}\;=\;\mathbf{x}_{t-1}\;+\;\Delta_{\theta}\!\bigl{(}\mathbf{x}_{t-1};\pi(t)\bigr{)}.$ (8)
where $\Delta_{\theta}$ predicts the point displacement conditioned on a temporal code $\pi(t)$. Such motion fields can be trained directly from 2D images through differentiable rendering, dispensing with explicit 3D supervision and delivering per-point motion estimates. However, this learning problem is normally ill-posed–especially under limited observations,so auxiliary information and extra regularization are often introduced for reasonable solution (Sec. 4.5).
@@ 3.1.4 Hybrid Motion
Most real-world scenes exhibit a mixture of motion types–rigid, articulated, and general non-rigid–whose simultaneous presence gives rise to hybrid motion. A prime example is the human body: a primarily articulated skeleton undergoes a global rigid transformation, while soft tissue, loose clothing, and hair introduce finer, highly non-rigid dynamics [77, 78]. Capturing this interplay demands a representation that is both structured enough to model global motion and flexible enough to accommodate local deviations.
@@
In this formulation, $\mathcal{T}_{\theta_{1}}$ represents a relatively constrained, interpretable global motion model (e.g., rigid transformation or articulated skeleton deformation), while $\Delta_{\theta_{2}}$ is typically implemented as a neural field that predicts residual displacements to capture complex, local deformations. Besides effective human and animal modeling, similar approaches have been applied to other domains with composite motion, including deformable objects with near-rigid parts and multi-object scenes with varying motion characteristics [72, 51]. This hierarchical decomposition offers several advantages: it maintains physical interpretability by isolating well-defined transformations; it reduces the complexity of the learning problem by having the neural component focus on residual details rather than the entire motion; and it provides explicit control over coarse motion while allowing the capture of fine details.
@@ 3.2.1 Representing via 4D Spacetime
Starting from coordinate-based representations for static scenes [79, 1], which utilize the 5D plenoptic function to represent radiance fields as $\mathcal{F}_{\theta}\colon(\mathbf{x},\mathbf{d})\longmapsto(\mathbf{c},\sigma)$, approaches for dynamic scenes naturally extend the input domain to 6D by incorporating temporal information $t$ [26, 80, 16, 40], resulting in $\mathcal{F}_{\theta}\colon(\mathbf{x},\mathbf{d},t)\longmapsto(\mathbf{c},\sigma)$. More generally, methods may condition on alternative temporal encodings $\pi(t)$ beyond direct time input, such as frame indices $i$ or learnable per-frame latent codes $\ell_{i}$ [81, 41, 82].
With this formulation, each frame of the dynamic scene is optimized independently through frame-by-frame optimization, leveraging dense observations to minimize a rendering loss between the synthesized output $I_{r}$ and the ground-truth image $I_{gt}$:
$\arg\min_{\theta}\sum\left\|I_{r}-I_{gt}\right\|_{2}^{2},\ \ I_{r}=\mathcal{R}\Bigl{(}\mathcal{F}_{\theta}\bigl{(}\mathbf{x},\mathbf{d},\pi(t)\bigr{)}\Bigr{)},$ (10)
where $\mathcal{R}$ represents the differentiable volume rendering or splatting function. Motion is thus implicitly encoded within the same radiance field that represents the scene–rather than being handled by a separate motion field–and is supervised solely through available 2D image observations.
3.2.2 Representing via Canonical Space
A prevalent approach for modeling dynamic scenes decomposes scenes into a static canonical space and time-varying deformation fields. This canonical space–often designated as a "reference frame"–serves as a common coordinate system from which all observed frames are derived through learned deformations. For rigid or articulated objects, this canonical space could be intuitively defined, e.g., initial state for rigid objects or neutral pose for articulated objects. For general scenes, it should capture sufficient geometric detail to facilitate robust correspondence estimation across the sequence.
@@ 3.2.2 Representing via Canonical Space
$\Delta_{i\to c}(\mathbf{x}_{i})\;=\;\Psi_{\theta}\bigl{(}\mathbf{x}_{i};\ \pi(t)\bigr{)}.$ (12)
@@
Within this framework, point correspondence between any two observation frames $i$ and $j$ could be set up by taking the shared canonical space as a intermediate station:
@@ 3.2.3 Representing via Flow Field
A more direct strategy for modeling dynamic scenes involves capturing motion between consecutive frames rather than relating observations to a shared reference. This frame-to-frame approach leverages incremental deformations to represent complex motions, decomposing significant transformations into smaller, more tractable steps. Such representations naturally accommodate topology changes and extreme deformations that challenge canonical space methods.
The flow field provides a principled framework for modeling these consecutive-frame dynamics. In their continuous form, velocity fields [85, 50, 19] specify instantaneous motion by assigning a velocity vector $\mathbf{v}(\mathbf{x},t)$ to each point $\mathbf{x}$ in space at time $t$. The point movement of point x from time $t-1$ to $t$ can be represented by integrating this velocity field:
@@
where $\mathbf{v}(\mathbf{x}(\tau),\tau)$ represents the velocity at intermediate time $\tau$. While theoretically elegant, obtaining continuous ground-truth velocities is often practically infeasible.
For discrete time steps, scene flow fields $\mathbf{O}(\mathbf{x},t)$ directly model displacement vectors between consecutive frames:
$\mathbf{O}(\mathbf{x},t)=\Delta_{t}\;=\;\mathbf{x}_{t}\;-\;\mathbf{x}_{t-1}.$ (15)
Scene flow and velocity fields are mathematically related–velocity represents the time derivative of scene flow–expressed differentially as:
$\frac{\partial\mathbf{O}}{\partial t}(\mathbf{x},t)\;=\;\mathbf{v}\bigl{(}\mathbf{O}(\mathbf{x},t),\,t\bigr{)},\quad\text{s.t.}\quad\mathbf{O}(\mathbf{x},\,t-1)=\mathbf{x}_{t-1}.$ (16)
In practical implementations, both fields are typically parameterized using neural networks as $\mathbf{v}_{\theta}$ and $\mathbf{O}_{\theta}$ and optimized with scene radiance field through differentiable rendering.
These local flow fields can be composed to represent extended movements between far-away frames through: